[figure: 9.7 cm]

Magnetic properties of SmMn_{12−x}Ni_{x} system

W. Suski^{a,b}, K. Wochowski^{b}, T. Mydlarz^{a} and B. Belan^{c}

aInternational Lab. of High Magnetic Fields and Low Temperatures, Wrocław, Poland

bPolish Academy of Sciences, W .Trzebiatowski Institute of Low Temperature and

Structure Research, ul. Okólna 2, 50–422 Wrocław, Poland

cDepartment of Inorganic Chemistry, Ivan Franko National University of L’viv, Ukraine

The SmMn_{12−x}Ni_{x} alloys crystallize in the tetragonal, ThMn_{12}-type structure and they exist for the Ni concentration 4 ≤ x ≤ 8. The magnetic susceptibility has been investigated at T = 1.9–300 K in magnetic field of 5 kOe using the SQUID magnetometer. At 4.2 K the magnetization has been measured in magnetic field up to 140 kOe on polycrystalline and finely powdered samples. The temperature dependence of the magnetic susceptibility χ exhibits for the samples with x = 8, 7, 5 and 4 a maximum at T_{N} (?) ∼ 70 K, whereas the χ(T) plot for the x = 6 sample shows two maxima: at T ∼ 30 K and very diffuse at T ∼ 180 K. The field dependence of magnetization at low temperature is in principle linear. These results suggest that the maximum at about 70 K corresponds to the onset of an antiferromagnetic ordering of the Sm sublattice because it is not sensitive to the composition change. The double maxima for the x = 6 alloy seems to result from some crystallographic imperfection of the sample. As it was shown for the RMn_{12−x}T_{x}, where R — rare earth and T = Fe or Co, the magnetic ordering in these sublattices as well as in the Mn one is very sensitive to mutual relation of T and Mn concentration. In the present system both the Mn and Ni sublattices seem to be nonmagnetic.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : W. Suski

Address for correspondence : Polish Academy of Sciences, W .Trzebiatowski Institute of Low Temperature and Structure Research, P.O. Box 1410, 50–950 Wrocław 2, Poland

Email address : suski@int.pan.wroc.pl